An amusement park ride consists of a rotating circular platform 12.2 m in diameter from which 10 kg seats are suspended at the end of 2.76 m massless chains. When the system rotates, the chains make an angle of 25.2 ◦ with the vertical. The acceleration of gravity is 9.8 m/s 2 . What is the speed of each seat? Answer in units of m/s. 



1
Expert's answer
2022-01-19T10:11:56-0500

"\tan\alpha=\frac{v^2}{gr},"

"v=\sqrt{{gr\tan\alpha}}=11~\frac ms."